# Effects of the entropy source on Monte Carlo simulations

### v1

@@ Abstract
In this paper we show how different sources of random numbers influence the outcomes of Monte Carlo simulations. We compare industry-standard pseudo-random number generators (PRNGs) to a quantum random number generator (QRNG) and show, using examples of Monte Carlo simulations with exact solutions, that the QRNG yields statistically significantly better approximations than the PRNGs. Our results demonstrate that higher accuracy can be achieved in the commonly known Monte Carlo method for approximating $\pi$. For Buffon’s needle experiment, we further quantify a potential reduction in approximation errors by up to 1.89×1.89$\times$1.89 × for optimal parameter choices when using a QRNG and a reduction of the sample size by 8$\sim$ 8$\times$ 8 × for sub-optimal parameter choices. We attribute the observed higher accuracy to the underlying differences in the random sampling, where a uniformity analysis reveals a tendency of the QRNG to sample the solution space more homogeneously. Additionally, we compare the results obtained with the QRNG and PRNG in solving the non-linear stochastic Schrödinger equation, benchmarked against the analytical solution. We observe higher accuracy of the approximations of the QRNG and demonstrate that equivalent results can be achieved at 1/3 to 1/10-th of the costs.
@@ 1 Introduction
Stochastic simulations have been one of the earliest applications of large-scale computer simulations in today’s science and industry. Monte Carlo (MC) simulations, which use random numbers as an integral part of obtaining a solution of stochastic problems, were first developed by Neumann and Ulam during the Manhattan project, and are used to this day in planning radiation therapy [1], protein folding [2], particle methods in general [3] and the modelling of societal segregation [4, 5]. Evolution of stochastic processes, described by stochastic differential equations (SDE), is a mainstay of the financial industry (such as the Black-Scholes equation), and underpins the development of optical systems [6, 7] as well as high-precision measurements [8]. The numerical solution of a stochastic differential equation is a specific case of a Monte Carlo method. Additionally, the growing adoption of uncertainty quantification, to include uncertainties in modelling as well as the stochastic nature of real-world observations, introduces Monte-Carlo methods to areas in which these have previously played only a niche role.
Fundamental to all of the fields listed above is the availability of random numbers on the platform used for the simulations. It is, however, commonly known that all ”random” numbers available in system libraries are simulated by algorithms generating deterministic sequences and are therefore not truly random, [9]. Mistakes in the past, such as highly correlated RANDU generator[10], have resulted in the development of high-quality pseudo-random number generators (PRNGs), of which the Mersenne Twister (MT) [11] is nowadays seen as the industry standard, being the standard engine adopted in Python [12] (and NumPy up to v. 1.15 [13]) and a common example introducing PRNGs in C++[14].
Scientific applications of Monte Carlo methods, relying on repeating the same operations, are especially popular due to the relative ease of their implementation and parallelisation, allowing for a simple utilisation of all available compute resources. However, the PRNGs described above are, in general, not suited for practical parallel applications. To fill this void, parallel PRNG libraries exist, such as Tina’s random number generator library or the scalable parallel random number generator library [15].
@@
In the following, we introduce models used to investigate the suitability of various random number sources for Monte Carlo applications and solutions to stochastic differential equations as well as the tests used to ascertain the significance of the observations.
@@ 2 Results
A MC simulation samples, in general, an unknown distribution thereby providing an empirical approximation to said distribution. We use the archetypal MC approximation of $\pi$ as the initial stochastic test bed for our analysis to assess how different sources of randomness affect the simulation outcomes. This choice is motivated by the availability of an exact solution, simplicity of the procedure as well as the model being the universal example for MC methods. We employ the perhaps best known $\tfrac{\pi}{4}$ approximation (Method A, c.f. fig. 1C), i.e. generating random points inside a unit square and counting the proportion that falls inside the quarter circle, as well as ”Buffon’s needle” approximation of $\tfrac{2}{\pi}$ (Method B, c.f fig. 1D)[10]. Due to the different quantities being approximated ($\tfrac{\pi}{4}\;\text{vs.}\;\tfrac{2}{\pi}$), we expect the uncertainties of the resulting $\pi$ approximations to behave differently with respect to the number of points $M$ on $[0,1]^{2}$ used to obtain one sample $\tilde{\pi}_{i}$ of $\pi$ as well as the number of samples $N$ used to obtain an approximation of $\bar{\pi}=\tfrac{1}{N}\sum_{i=1}^{N}\tilde{\pi}_{i}$. Comparing a self-certifying quantum random number generator (c.f. fig. 1A and Supplementary Materials SM A.1.2) to industry-standard PRNGs (SM A.1.1), we demonstrate that the QRNG leads to better approximations than the PRNGs for both methods. We show that the results obtained with the QRNG pass the sign test and t-tests for larger sample sizes than the PRNG: on the one hand, this permits a reduction in sample size while maintaining the same accuracy as achieved by the PRNG. On the other hand, a better approximation of $\pi$ is achieved when considering the same number of samples.
To investigate the observed differences further, we then perform an uniformity analysis to assess underlying sampling itself, before moving to an applied example comparing the QRNG to a PRNG in solving the stochastic Schrödinger equation (SSE).
@@
First we consider the commonly known method of estimating $\pi$ (Method A, c.f. SM sec. A.2) and apply the sign test to $N=2.832\cdot 10^{6}$ $\tilde{\pi}$ samples (c.f. SM sec. A.4), each obtained using $M=10^{5}$ points on $[0,1]^{2}$. The coordinates $(x,y)$ of each point are drawn from the uniform distribution between 0 and 1 using a QRNG or Mersenne Twister PRNG, respectively. We consider the $99.99\%$ confidence interval of the hypothesis $H_{0}:m=\pi$ (i.e. the median is $\pi$), with the alternative hypothesis $H_{A}:m\neq\pi$. Here, fig. 1 E illustrates the evolution of the acceptance interval (shown as the green bar) for growing sample sizes, along with the values of the test statistic (number of values $>m$) for the datasets obtained with PRNG (blue) and QRNG (red) and the ideal statistic ($\frac{N}{2}$, gold). The following can be observed: after roughly $N_{C}\sim 1.4\cdot 10^{6}$ $\tilde{\pi}$ samples, $H_{0}$ must be rejected for the data obtained using the Mersenne Twister PRNG (i.e. the test statistic leaves the acceptance range) but it can not be rejected for data obtained using the QRNG. Explicitly this means that we must reject the claim that the median is $\pi$ for PRNG data but do not reject it for QRNG data: the latter may thus be considered to provide qualitatively better results. We emphasise that no statement is made about a statistical error of the 2nd kind ($H_{A}$ is true but $H_{0}$ is accepted) for the QRNG dataset, i.e., the analysis does not bound the probability of $H_{A}:m\neq\pi$ for the QRNG dataset.
@@
Since Buffon’s method (method B) approximates $\frac{2}{\pi}$ instead of $\frac{\pi}{4}$, the behaviour of the $\tilde{\pi}$ distributions is different. In the Supplementary Materials fig. 8, we show that the mean obtained with method B differs from the true value of $\pi$ more than the mean obtained with method A, indicating that method A is the better choice for approximating $\pi$ via MC simulations. However, although method B yields larger approximation errors for $M\leq 5\cdot 10^{5}$, it displays a clear dependence on the number of points-per-sample (pps) $M$ by following the expected $\frac{1}{\sqrt{M}}$ convergence, resulting in a better approximation at $M=10^{6}$ pps and beyond (c.f. SM fig. 9. In conjunction with the analysis presented below, we conclude that method B allows for a clear distinction of the effects of the sample size $N$ and the points-per-sample $M$, whereas for method A such a distinction is not easily made.
@@
The funnel-shaped shaded blue region corresponds to the $95\%$ acceptance range obtained by reordering the $t$-test equation (c.f. SM 8) with the upper and lower bounds obtained utilising the empirical variance of the joint pPRNG and QRNG dataset. Additionally, multi-coloured regions of the x-axis indicate within which execution of the code the data has been obtained, each of which corresponds to a pPRNG initialisation with a different seed (c.f. SM sec. B). We observe in fig. 2b that the pPRNG mean is consistently rejected as an approximation of $\pi$ after $N=2\cdot 10^{5}$ samples, whereas the QRNG mean is rejected only after $N=10^{6}$ samples. Here we must note that the values converge to the true value only in the case $M\rightarrow\infty$, while, for a finite $M$, the limit value is unknown but larger than $\pi$. We have devised an estimator (c.f. SM eq. 2) for the difference between the limit value and $\pi$ at a given pps $M$ in SM sec. A.2. The expected mean is shown as the black dotted line in the figure SM fig. 9 and is in good agreement ($\mu_{data}-\mu_{expect}=3.8\cdot 10^{-6}$) with the empirical mean taken across both datasets.
@@
Here, evidently, the approximation error (normalised deviation) when using Buffon’s method decreases with growing number of samples $N$, as predicted by the central limit theorem. However, after $N_{PRNG}^{C}\approx 10^{5}$ samples obtained with the pPRNG ($N_{QRNG}^{C}\approx 3\cdot 10^{5}$ for QRNG), the behaviour changes drastically, converging to a limit value $\tilde{\epsilon}\approx 2\cdot 10^{-3}$ (c.f. SM fig. 9).
We denote the largest sample size after which $\tilde{\epsilon}_{N}$ exceeds the $1/\sqrt{N}$ threshold as $N^{C}_{QRNG,PRNG}$. Beyond this threshold $\tilde{\epsilon}_{N}$ breaks the expected convergence behaviour and t-test acceptance bounds (indicated by the dashed lines in the plot), thus resulting in no further increase in accuracy. We note that $N^{C}_{PRNG}\approx M$ for various $M$ and, in view of SM fig. 9 interpret that the points-per-sample $M$ dictate the ”trueness” of the approximation, whereas the number of samples $N$ determines the precision of the measurement (c.f SM fig. 10). For $N\leq M$ the accuracy is dictated by the trueness of the observation, whereas for $N>M$ it is determined by the precision of each observation. This implies that the number of points-per-sample $M$ must scale as $\sim N$ to maintain the $\sim 1/\sqrt{N}$ convergence. Hence a simulation with $N>N^{C}\approx M$ is of little use as the approximation error plateaus for $N>M$ (c.f. fig. 2c). The same holds true for method A, albeit there the distinction between $M$ and $N$ is less clear as both affect the width of the distribution.
@@
As shown in SM fig. 8, the samples are approximately normally distributed and as such use of the t-test is appropriate. To obtain the $95\%$ acceptance threshold of the t-test, the $\frac{1}{\sqrt{N}}$ curve must be scaled by $2.05$ (c.f. SM 8) resulting in the green dashed line in fig. 2c. The range of acceptable deviations is thus larger: when we apply the same analysis as above, this results in $N_{PRNG}^{C}=2.3\cdot 10^{5}=2.3M$ and $N_{QRNG}^{C}=10^{6}=10M$. Thus using the QRNG results in a 4.54$\sim$ 4.54$\times$ 4.54 × larger admissible sample size and an approximation error which is 1.89$\sim$ 1.89$\times$ 1.89 × smaller than that of the pPRNG approximation.
@@
To better understand the observations made with the $\pi$ model, we consider the underlying sampling, or more specifically, the uniformity of the point distribution on $[0,1]^{2}$ as provided by the QRNG and the Mersenne Twister[11] PRNG. Uniformity was analysed using the nearest-neighbour distance ($N\!N$), largest empty sphere ($LES$), dispersion ($C_{V}$, $C\!E$) and potential energy metrics ($Q$ and $R_{nr}$)(SM sec. A.3). We remark that an analysis of the uniformity of sampling of a domain requires an analysis of the relation of the point distribution to the boundary. In absence of such the homogeneity of the distribution is considered. Hence, in the following we refer to solely the behaviour of a point distribution, where $N\!N$ and $LES$ assess the distance between pairs of points, $C\!E$ and $R_{nr}$ test for clustering and ’non-randomness’, and $C_{V}$ and $Q$ provide more global measures for uniformly dispersion by averaging the contributions from each point. These measures, further described in SM sec. A.3, are collected 50 times for the sample points sets of various sizes and numerical precision both for sequential pre-generated QRNG and PRNG (Mersenne Twister). The 50 averages, each taken from 100 experiments with a given number of point samples, are summarised in the box plots.
@@
In fig. 3 we present results for the $50\times 100$ sets of 1000 points, each of which is defined by a single-precision number pair representing the $x$- and $y$-coordinates. Although the direct application of $N\!N$ measure does not illustrate any statistically significant difference, the measures derived from $N\!N$ lead to the following observations: the means and medians of quality measure $Q$ and $R_{nr}$ are significantly higher for QRNG as compared to PRNG but remain below 1, suggesting that QRNG offer a better dispersion of values, indicating a tendency towards more randomness. The observation that the Coefficient of Variation ($C_{V}$) is smaller for QRNG than for PRNG, despite no statistically significant difference observed in $N\!N$ distances, fig. 3, indicates that while both types of generators produce uniformly distributed numbers, QRNG does so with consistency and less variability more closely matching theoretical expectations. The statistically significant differences are observed at a $95\%$ confidence level between the sets of single-precision floating-point numbers generated by QRNG and PRNG, as determined by the Wilcoxon, Student’s T-test, Anderson-Darling, and Kolmogorov-Smirnov tests, and shown in the SM tbl. 1.

### v2

@@ Abstract
In this paper we show how different sources of random numbers influence the outcomes of Monte Carlo simulations. We compare industry-standard pseudo-random number generators (PRNGs) to a quantum random number generator (QRNG) and show, using examples of Monte Carlo simulations with exact solutions, that the QRNG yields statistically significantly better approximations than the PRNGs. Our results demonstrate that higher accuracy can be achieved in the commonly known Monte Carlo method for approximating $\pi$. For Buffon’s needle experiment, we further quantify a potential reduction in approximation errors by up to 1.89×1.89$\times$1.89 × for optimal parameter choices when using a QRNG and a reduction of the sample size by 8$\sim$ 8$\times$ 8 × for sub-optimal parameter choices. We attribute the observed higher accuracy to the underlying differences in the random sampling, where a uniformity analysis reveals a tendency of the QRNG to sample the solution space more homogeneously.
@@ 1 Introduction
Stochastic simulations have been one of the earliest applications of large-scale computer simulations in today’s science and industry. Monte Carlo (MC) simulations, which use random numbers as an integral part of obtaining a solution of stochastic problems, were first developed by Neumann and Ulam during the Manhattan project, and are used to this day in planning radiation therapy [1], protein folding [2], particle methods in general [3] and the modelling of societal segregation [4, 5]. Evolution of stochastic processes, described by stochastic differential equations (SDE), is a mainstay of the financial industry (such as the Black-Scholes equation), and underpins the development of optical systems [6, 7] as well as high-precision measurements [8]. The numerical solution of a stochastic differential equation is a specific case of a Monte Carlo method. Additionally, the growing adoption of uncertainty quantification, to include uncertainties in modelling as well as the stochastic nature of real-world observations, introduces Monte Carlo methods to areas in which these have previously played only a niche role.
Fundamental to all of the fields listed above is the availability of random numbers on the platform used for the simulations. It is, however, commonly known that all ”random” numbers available in system libraries are simulated by algorithms generating deterministic sequences and are therefore not truly random, [9]. Mistakes in the past, such as highly correlated RANDU generator[10], have resulted in the development of high-quality pseudo-random number generators (PRNGs), of which the Mersenne Twister (MT) [11] is nowadays seen as the industry standard, being the standard engine adopted in Python [12] (and NumPy up to v. 1.15 [13]) and a common example introducing PRNGs in C++[14]. Scientific applications of Monte Carlo methods, relying on repeating the same operations, are especially popular due to the relative ease of their implementation and parallelisation, allowing for a simple utilisation of all available compute resources. However, the PRNGs described above are, in general, not suited for practical parallel applications. To fill this void, parallel PRNG libraries exist, such as Tina’s random number generator library or the scalable parallel random number generator library [15].
@@
Even though reference examples, where benefits of QRNGs in MC simulations have been demonstrated, exist[21, 20], one might argue that benchmarking against PRNGs exhibiting periods shorter than the period of the ’minimum standard’ adopted in C++ [22] overemphasises the benefits of QRNGs. In principle, as has been demonstrated by Cirauqui et al. in [21], problems sensitive to correlations are suitable for an analysis of RNG quality. In the case delineated therein, the observables are computed from spins which, in turn, are directly dependent on the sequence of random numbers provided and thus an effective proxy measure for correlations within the RN sequence.
In this paper, we show that a QRNG will yield a better approximation than an industry-standard PRNG for a much simpler and cheaper (pertaining to the run time) model.
In the following, we introduce models used to investigate the suitability of various random number sources for Monte Carlo applications as well as the tests used to ascertain the significance of the observations.
@@ 2 Results
A MC simulation samples, in general, an unknown distribution thereby providing an empirical approximation to said distribution. We use the archetypal MC approximation of $\pi$ as the initial stochastic test bed for our analysis to assess how different sources of randomness affect the simulation outcomes. This choice is motivated by the availability of an exact solution, simplicity of the procedure as well as the model being the universal example for MC methods. We employ the perhaps best known $\tfrac{\pi}{4}$ approximation (Method A, c.f. fig. 1C), i.e. generating random points inside a unit square and counting the proportion that falls inside the quarter circle, as well as ”Buffon’s needle” approximation of $\tfrac{2}{\pi}$ (Method B, c.f fig. 1D)[23]. Due to the different quantities being approximated ($\tfrac{\pi}{4}\;\text{vs.}\;\tfrac{2}{\pi}$), we expect the uncertainties of the resulting $\pi$ approximations to behave differently with respect to the number of points $M$ on $[0,1]^{2}$ used to obtain one sample $\tilde{\pi}_{i}$ of $\pi$ as well as the number of samples $N$ used to obtain an approximation of $\bar{\pi}=\tfrac{1}{N}\sum_{i=1}^{N}\tilde{\pi}_{i}$. Comparing a self-certifying quantum random number generator (c.f. fig. 1A and Supplementary Materials SM sec. A.1.2) to industry-standard PRNGs (SM sec. A.1.1), we demonstrate that the QRNG leads to better approximations than the PRNGs for both methods. We show that the results obtained with the QRNG pass the sign test and t-tests for larger sample sizes than the PRNG: on the one hand, this permits a reduction in sample size while maintaining the same accuracy as achieved by the PRNG. On the other hand, a better approximation of $\pi$ is achieved when considering the same number of samples.
To investigate the observed differences further, we then perform an uniformity analysis to assess underlying sampling itself.
@@
First we consider the commonly known method of estimating $\pi$ (Method A, c.f. SM sec. A.2) and apply the sign test to $N=2.832\cdot 10^{6}$ $\tilde{\pi}$ samples (c.f. SM sec. A.4), each obtained using $M=10^{5}$ points on $[0,1]^{2}$. The coordinates $(x,y)$ of each point are drawn from the uniform distribution between 0 and 1 using a QRNG or Mersenne Twister PRNG, respectively. We consider the $99.99\%$ confidence interval of the hypothesis $H_{0}:m=\pi$ (i.e. the median is $\pi$), with the alternative hypothesis $H_{A}:m\neq\pi$. Here, fig. 1 E illustrates the evolution of the acceptance interval (shown as the green bar) for growing sample sizes, along with the values of the test statistic (number of values $>m$) for the datasets obtained with PRNG (blue) and QRNG (red) and the ideal statistic ($\frac{N}{2}$, gold). The following can be observed: after roughly $N_{C}\sim 1.4\cdot 10^{6}$ $\tilde{\pi}$ samples, $H_{0}$ must be rejected for the data obtained using the Mersenne Twister PRNG (i.e. the test statistic leaves the acceptance range) but it can not be rejected for data obtained using the QRNG. Explicitly this means that we must reject the claim that the median is $\pi$ for PRNG data but do not reject it for QRNG data: the latter may thus be considered to provide qualitatively better results
We emphasise that no statement is made about a statistical error of the 2nd kind ($H_{A}$ is true but $H_{0}$ is accepted) for the QRNG dataset, i.e., the analysis does not bound the probability of $H_{A}:m\neq\pi$ for the QRNG dataset.
@@
Since Buffon’s method (method B) approximates $\frac{2}{\pi}$ instead of $\frac{\pi}{4}$, the behaviour of the $\tilde{\pi}$ distributions is different. In the Supplementary Materials fig. 8, we show that the mean obtained with method B differs from the true value of $\pi$ more than the mean obtained with method A, indicating that method A is the better choice for approximating $\pi$ via MC simulations. However, although method B yields larger approximation errors for $M\leq 5\cdot 10^{5}$, it displays a clear dependence on the number of points-per-sample (pps) $M$ by following the expected $\frac{1}{\sqrt{M}}$ convergence, resulting in a better approximation at $M=10^{6}$ pps and beyond (c.f. fig. 2). In conjunction with the analysis presented below, we conclude that method B allows for a clear distinction of the effects of the sample size $N$ and the points-per-sample $M$, whereas for method A such a distinction is not easily made.
@@
The funnel-shaped shaded blue region corresponds to the $95\%$ acceptance range obtained by reordering the $t$-test equation (c.f. SM eq. 8) with the upper and lower bounds obtained utilising the empirical variance of the joint pPRNG and QRNG dataset. Additionally, multi-coloured regions of the x-axis indicate within which execution of the code the data has been obtained, each of which corresponds to a pPRNG initialisation with a different seed (c.f. SM sec. B). We observe in fig. 3b that the pPRNG mean is consistently rejected as an approximation of $\pi$ after $N=2\cdot 10^{5}$ samples, whereas the QRNG mean is rejected only after $N=10^{6}$ samples. Here we must note that the values converge to the true value only in the case $M\rightarrow\infty$, while, for a finite $M$, the limit value is unknown but larger than $\pi$. We have devised an estimator (c.f. SM eq. 2) for the difference between the limit value and $\pi$ at a given pps $M$ in SM sec. A.2. The expected mean is shown as the black dotted line in fig. 2 and is in good agreement ($\mu_{data}-\mu_{expect}=3.8\cdot 10^{-6}$) with the empirical mean taken across both datasets.
@@
Here, evidently, the approximation error (normalised deviation) when using Buffon’s method decreases with growing number of samples $N$, as predicted by the central limit theorem. However, after $N_{PRNG}^{C}\approx 10^{5}$ samples obtained with the pPRNG ($N_{QRNG}^{C}\approx 3\cdot 10^{5}$ for QRNG), the behaviour changes drastically, converging to a limit value $\tilde{\epsilon}\approx 2\cdot 10^{-3}$ (c.f. fig. 2).
We denote the largest sample size after which $\tilde{\epsilon}_{N}$ exceeds the $1/\sqrt{N}$ threshold as $N^{C}_{QRNG,PRNG}$. Beyond this threshold $\tilde{\epsilon}_{N}$ breaks the expected convergence behaviour and t-test acceptance bounds (indicated by the dashed lines in the plot), thus resulting in no further increase in accuracy. We note that $N^{C}_{PRNG}\approx M$ for various $M$ and, in view of fig. 2, interpret that the points-per-sample $M$ dictate the ”trueness” of the approximation, whereas the number of samples $N$ determines the precision of the measurement (c.f SM fig. 9). For $N\leq M$ the accuracy is dictated by the trueness of the observation, whereas for $N>M$ it is determined by the precision of each observation. This implies that the number of points-per-sample $M$ must scale as $\sim N$ to maintain the $\sim 1/\sqrt{N}$ convergence. Hence, a simulation with $N>N^{C}\approx M$ is of little use as the approximation error plateaus for $N>M$ (c.f. fig. 3c). The same holds true for method A, albeit there the distinction between $M$ and $N$ is less clear as both affect the width of the distribution.
@@
As shown in SM fig. 8, the samples are approximately normally distributed and as such use of the t-test is appropriate. To obtain the $95\%$ acceptance threshold of the t-test, the $\frac{1}{\sqrt{N}}$ curve must be scaled by $2.05$ (c.f. SM eq. 8) resulting in the green dashed line in fig. 3c. The range of acceptable deviations is thus larger: when we apply the same analysis as above, this results in $N_{PRNG}^{C}=2.3\cdot 10^{5}=2.3M$ and $N_{QRNG}^{C}=10^{6}=10M$. Thus using the QRNG results in a 4.54$\sim$ 4.54$\times$ 4.54 × larger admissible sample size and an approximation error which is 1.89$\sim$ 1.89$\times$ 1.89 × smaller than that of the pPRNG approximation.
@@
Figure 4: Deviation (error) of the computed approximation from $\pi$ at a fixed pps $M$ for data obtained with the QRNG and two different PRNGs.
In fig. 4, we augment fig. 3c to contrast the behaviour of the QRNG (blue) and pPRNG (red), exhibiting a large period, to that of the ”Minimum standard” linear congruential generator[22] (LCG, black) as it is implemented in the C++11 standard. First note that the LCG dataset size was capped at $1/3$ of the pPRNG dataset size since the approximation error does not decrease beyond $\sim N_{QRNG}^{C}$. The normalised deviation $\tilde{\epsilon}_{N}$ displays several notable features for the LCG that are not present in either the QRNG or pPRNG equivalents. For $N>N_{QRNG}^{C}$, it reaches, on average, an asymptotic value of $\tilde{\epsilon}_{LCG}=1.24\cdot 10^{-2}$, or 6.12$\approx$ 6.12$\times$ 6.12 × larger than the deviation obtained with QRNG data.
Further we observe that $N_{LCG}^{C}\approx 1.9\cdot 10^{4}$ which is 12$\sim$ 12$\times$ 12 × smaller than $N_{PRNG}^{C}$. Finally, $\tilde{\epsilon}_{LCG}$ displays a periodic structure for $N\gtrsim N_{LCG}^{C}$, whose oscillation magnitude decreases, but does not vanish, with increasing sample size. The oscillation period is directly related to the period $m=2^{31}-1\approxeq 2.15\cdot 10^{9}$ of the LCG, which, for the given $M$, translates to a period $N\approxeq 10737$ of the $\tilde{\pi}$. Given the notably higher approximation error $\tilde{\epsilon}$, $N_{LCG}^{C}$ being an order of magnitude lower than $N_{PRNG}^{C}$, and the obvious effects of the LCG period, we have excluded this LCG from our experiments and, in contrast to other works [21], not considered PRNGs of shorter periods than the ”Minimum standard”.
@@
To better understand the observations made with the $\pi$ models, we consider the underlying sampling, or more specifically, the uniformity of the point distribution on $[0,1]^{2}$ as provided by the QRNG and the Mersenne Twister[11] PRNG. Uniformity was analysed using the nearest-neighbour distance ($N\!N$), largest empty sphere ($LES$), dispersion ($C_{V}$, $C\!E$) and potential energy metrics ($Q$ and $R_{nr}$)(SM sec. A.3). We remark that an analysis of the uniformity of sampling of a domain requires an analysis of the relation of the point distribution to the boundary. In absence of such the homogeneity of the distribution is considered. Hence, in the following we refer to solely the behaviour of a point distribution, where $N\!N$ and $LES$ assess the distance between pairs of points, $C\!E$ and $R_{nr}$ test for clustering and ’non-randomness’, and $C_{V}$ and $Q$ provide more global measures for uniformly dispersion by averaging the contributions from each point. These measures, further described in SM sec. A.3, are collected 50 times for the sample points sets of various sizes and numerical precision both for sequential pre-generated QRNG and PRNG (Mersenne Twister). The 50 averages, each taken from 100 experiments with a given number of point samples, are summarised in the box plots.
@@
W test T test AD test KS test
$N\!N$ 0.5210 0.5286 0.2500 0.8693
$Q$ 0.0171 0.0141 0.0155 0.0217
$C_{V}$ 0.5148 0.5577 0.2500 0.9977
$C\!E$ 0.0207 0.0156 0.0183 0.0392
$R_{nr}$ 0.0207 0.0156 0.0183 0.0392
$LES$ 0.6390 0.6575 0.2500 0.5487
Table 1: $p$-values for 1000 float points, $\alpha=0.05$
In fig. 5 we present results for the $50\times 100$ sets of 1000 points, each of which is defined by a single-precision number pair representing the $x$- and $y$-coordinates. Although the direct application of $N\!N$ measure does not illustrate any statistically significant difference, the measures derived from $N\!N$ lead to the following observations: the means and medians of the quality measure $Q$ and $R_{nr}$ are significantly higher for QRNG as compared to PRNG but remain below 1, suggesting that QRNG offer a better dispersion of values, indicating a tendency towards more randomness. The observation that the Coefficient of Variation ($C_{V}$) is smaller for QRNG than for PRNG, despite no statistically significant difference observed in $N\!N$ distances, fig. 5, indicates that while both types of generators produce uniformly distributed numbers, QRNG does so with consistency and less variability more closely matching theoretical expectations. The statistically significant differences are observed at a $95\%$ confidence level between the sets of single-precision floating-point numbers generated by QRNG and PRNG, as determined by the Wilcoxon, Student’s T-test, Anderson-Darling, and Kolmogorov-Smirnov tests, and shown in tbl. 1.